# Implicit Style-Content Separation using B-LoRA

### v1

@@ Appendix 0.A Comparisons
In Section 5 of the main paper, we conducted a comparison of our B-LoRA method against four state-of-the-art baselines for image stylization incorporating personalization [cnetSDXL, ziplora23, sohn2023styledrop, styleHertz23]. In this section, we delve deeper into the implementation details and present additional qualitative results. To begin, we employed DreamBooth-LoRA [DBlora] fine-tuning to obtain both style and content LoRAs, utilizing the same parameter configuration as ZipLoRA [ziplora23]. For content images, we conducted fine-tuning across a set of images of the same object, except for the experiment involving a single image. However, for style LoRAs, we conducted fine-tuning using a single style image. We utilized the prompts provided in DreamBooth [dreambooth23] and StyleDrop [sohn2023styledrop], specifically “A [v] <object>” or “A <object> in [s] style” for content and style, respectively. Subsequently, for ControlNet combined with DreamBooth-LoRA, we leveraged the publicly available implementation of ControlNet on SDXL from huggingface [cnetSDXL]. this approach involved utilizing the style LoRAs we trained for style transfer while employing CannyEdge with thresholds of 100 and 200 for content guidance in ControlNet. For StyleDrop [sohn2023styledrop], we followed the methodology outlined in StyleAligned [styleHertz23] for fine-tuning the model over the style images, followed by fusing the content LoRAs with the SDXL weights. Similarly, for StyleAligned [styleHertz23], we utilized the authors’ implementation for subject-driven generation alongside our content LoRAs. Lastly, for ZipLoRA [ziplora23], we use the unofficial implementation [unofficialZipLoRA] with default parameters. We provide additional comparisons of our B-LoRA method against the aforementioned approaches using the same evaluation set presented in Section 5 of the main paper. These additional comparisons are illustrated in Figure 2. Furthermore, we provide comparisons with challenging content inputs, such as stylized images, presented in Figure 3. We also showcase comparisons with challenging style inputs, such as object images, in Figure 4. These examples demonstrate the robustness of our method in handling diverse and complex content and style references.
@@
Appendix 0.B Limitations
@@
Appendix 0.C Analysis and Ablation
@@
As detailed in Section 4 of the main paper, the SDXL UNet comprises 11 transformer blocks, with the high-resolution blocks containing 2 attention layers each and the middle 6 blocks containing 10 attention layers each (see Figure 3 in the main paper). To explore the impact of different block combinations on the resulting image, we divided the UNet into 8 blocks $\{W^{0}_{0}\ldots W^{7}_{0}\}$, where $\{W^{1}_{0}\ldots W^{6}_{0}\}$ represent the bottleneck blocks, as discussed in Section 4, and designated $W^{0}_{0}$ and $W^{7}_{0}$ for the high-resolution blocks at the edges. We aimed to assess the effects of optimizing various block combinations $\{\Delta W^{i},\Delta W^{j}\}$ by jointly training the LoRA weights of the corresponding blocks. Qualitative results are depicted in Figures 7 and 8, where each cell (i, j) represents the reconstruction image for the prompt “A [v]” after training the LoRAs solely for the i-th and j-th blocks of the SDXL Unet. The diagonal entries represent output generated by training a single block. Upon examination, we observed that optimizing $\{\Delta W^{4},\Delta W^{5}\}$ consistently produced the most satisfactory results for content and style, respectively, outperforming other combinations. Notably, the reconstruction in cell (4, 5) yielded the best results achievable among all combinations, supporting our findings in the main paper. Furthermore, we noted that the combination of blocks 2 and 5 also achieved satisfactory reconstruction. However, employing this combination may lead to less disentanglement of style from content, as $\Delta W^{5}$ needs to “cover” $\Delta W^{2}$ by learning content details instead of focusing primarily on style, as intended. This observation further solidifies our choice of optimizing $\{\Delta W^{4},\Delta W^{5}\}$ for effective style-content separation.
@@
To validate our choice of the prompt “A [v]” during optimization, we conducted an ablation study regarding the prompt used during training. As described in the DreamBooth [dreambooth23] paper, the authors suggest that the most efficient way to conduct the fine-tuning process is by using the prompt “A [v] <class-name>”, where [v] is the token dedicated for personalization, and <class-name> is the class of the object depicted in the input image. We compare our method of optimizing $\Delta W^{4}$ and $\Delta W^{5}$ with the prompt “A [v]” against using the suggested “A [v] <class-name>” prompt.
@@
Appendix 0.D B-LoRA for Personalization
Throughout the paper, our method has been implemented using a single image for decoupling style and content. However, by training our method using multiple images for content, we can recontextualize reference objects while preserving stylization quality. In Figures 10 and 11, we showcase the versatility of our method by combining various B-LoRAs for style and content with text prompts. Note that the style can be derived from the reference style or from other objects.
@@ Appendix 0.D B-LoRA for Personalization
Appendix 0.E Additional Results
Our B-LoRA method focuses on three main applications: image stylization based on image style references, text-based image stylization, and consistent style generation. In Figures 12 and 13, we present additional results generated by our approach for image stylization based on image style references. The columns represent the style reference images, while the rows correspond to the content reference images. As discussed, our method demonstrates proficiency in extracting content from style images (Figure 14) and extracting style from objects for object mixing tasks (Figure 15). In Appendices 0.E and 0.E, we provide qualitative results showcasing our method’s performance on randomly selected objects and styles from our evaluation set. These examples further highlight the robustness of our approach to handling diverse content and style references. In Figure 18 we present additional qualitative results for text-based image stylization. As discussed in the paper, by utilizing only the learned B-LoRA weights capturing the content, our method enables text-guided style manipulation while effectively preserving the input object’s content and structure. These results demonstrate the flexibility of our approach in allowing challenging style manipulations through textual guidance.
@@ Appendix 0.E Additional Results
ContentStyle &   Refer to caption Refer to caption Refer to caption Refer to caption
@@
c | @   c c c c c c
Input
ContentStyle &   Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
c | @   c c c c c c

### v2

@@ Appendix A Comparisons
In Section 5 of the main paper, we conducted a comparison of our B-LoRA method against four state-of-the-art baselines for image stylization incorporating personalization [27, 47, 48, 22]. In this section, we delve deeper into the implementation details and present additional qualitative results. To begin, we employed DreamBooth-LoRA [46] fine-tuning to obtain both style and content LoRAs, utilizing the same parameter configuration as ZipLoRA [47]. For content images, we conducted fine-tuning across a set of images of the same object, except for the experiment involving a single image. However, for style LoRAs, we conducted fine-tuning using a single style image. We utilized the prompts provided in DreamBooth [45] and StyleDrop [48], specifically “A [v] <object>” or “A <object> in [s] style” for content and style, respectively. Subsequently, for ControlNet combined with DreamBooth-LoRA, we leveraged the publicly available implementation of ControlNet on SDXL from huggingface [27]. this approach involved utilizing the style LoRAs we trained for style transfer while employing CannyEdge with thresholds of 100 and 200 for content guidance in ControlNet. For StyleDrop [48], we followed the methodology outlined in StyleAligned [22] for fine-tuning the model over the style images, followed by fusing the content LoRAs with the SDXL weights. Similarly, for StyleAligned [22], we utilized the authors’ implementation for subject-driven generation alongside our content LoRAs. Lastly, for ZipLoRA [47], we use the unofficial implementation [36] with default parameters. We provide additional comparisons of our B-LoRA method against the aforementioned approaches using the same evaluation set presented in Section 5 of the main paper. These additional comparisons are illustrated in Figure 11. Furthermore, we provide comparisons with challenging content inputs, such as stylized images, presented in Figure 12. We also showcase comparisons with challenging style inputs, such as object images, in Figure 13. These examples demonstrate the robustness of our method in handling diverse and complex content and style references.
@@
Comparisons to Baselines Beyond SDXL-Based Approaches
We provide additional comparisons of our method with three other image stylization techniques that do not rely on SDXL: StyTr2 [12], AdaAttn [35], and SWAG [55]. We evaluated the results using the same quantitative metrics described in the main paper. Figure 14 presents a qualitative comparison of the same set shown in the main paper, and Table 2 contains the quantitative results.
Table 2: Quantitative comparison: We measure the average cosine similarity between the DINO features of the output image and the reference style and content. In this experiment, we use a single input image for evaluation.
StyTr2 AdaAttn SWAG Ours
Style
Transfer
$0.83$ $0.818$ $0.883$ $0.881$
Content $0.854$ $0.828$ $0.788$ $0.790$
Content Style StyTr2 AdaAttn SWAG Ours
Figure 14: Comparison with alternative approaches that do not rely on SDXL. The input style and content references are shown on the left.
Comparisons to InstantStyle
InstantStyle [54] is a concurrent work to ours. Aimed at performing image stylization tasks based on a style image reference. InstantStyle achieves this by injecting the CLIP embedding of the style image into style-specific blocks within SDXL, similar to our method, where the fifth block is selected for the style condition. Notably, InstantStyle uses a trained IP-Adapter model and does not require fine-tuning, which is its main advantage over our method. Both approaches provide compelling results in consistent style generation, as presented in Figure 15. For content conditioning, InstantStyle utilizes ControlNet, while our method separates content from style and extracts both. This allows for better content preservation in scenarios where ControlNet may not capture the content well enough or may override the style, as shown in Figure 16. Additionally, InstantStyle requires the content component to be explicitly defined to subtract its CLIP embedding from the style embedding, whereas our approach learns the content and style implicitly. For a fair comparison, we trained our method using the style images from InstantStyle.
Input Style Method “A dog” “A moose” “A giraffe” “A girl”
Refer to caption Ours Refer to caption Refer to caption Refer to caption Refer to caption
InstantStyle Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Ours Refer to caption Refer to caption Refer to caption Refer to caption
InstantStyle Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Ours Refer to caption Refer to caption Refer to caption Refer to caption
InstantStyle Refer to caption Refer to caption Refer to caption Refer to caption
Figure 15: Comparison of stylization results between our method and InstantStyle. The input style image is shown in the first column, followed by results generated by our method and InstantStyle for different prompts: “A dog”, “A moose”, “A giraffe”, and “A girl”. The images of InstantStyle are taken from the original paper. Both approaches achieve consistent style generation, demonstrating the effectiveness of style transfer.
Content Style InstantStyle Ours
Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption
Figure 16: Comparison of style and content mixing between our method and InstantStyle. The results illustrate cases where ControlNet, used by InstantStyle, may fail to adequately capture the content or may override the style. For example, in the fourth row, we can see that ControlNet failed to extract the shape of the dog, leading to unsatisfactory results, While our method demonstrates better content preservation. The images showcase the stylization applied to various content images, highlighting differences in how each approach handles content and style integration.
Appendix B Limitations
@@
Appendix C Analysis and Ablation
@@
As detailed in Section 4 of the main paper, the SDXL UNet comprises 11 transformer blocks, with the high-resolution blocks containing 2 attention layers each and the middle 6 blocks containing 10 attention layers each (see Figure 3 in the main paper). To explore the impact of different block combinations on the resulting image, we divided the UNet into 8 blocks $\{W^{0}_{0}\ldots W^{7}_{0}\}$, where $\{W^{1}_{0}\ldots W^{6}_{0}\}$ represent the bottleneck blocks, as discussed in Section 4, and designated $W^{0}_{0}$ and $W^{7}_{0}$ for the high-resolution blocks at the edges. We aimed to assess the effects of optimizing various block combinations $\{\Delta W^{i},\Delta W^{j}\}$ by jointly training the LoRA weights of the corresponding blocks. Qualitative results are depicted in Figures 19 and 20, where each cell (i, j) represents the reconstruction image for the prompt “A [v]” after training the LoRAs solely for the i-th and j-th blocks of the SDXL Unet. The diagonal entries represent output generated by training a single block. Upon examination, we observed that optimizing $\{\Delta W^{4},\Delta W^{5}\}$ consistently produced the most satisfactory results for content and style, respectively, outperforming other combinations. Notably, the reconstruction in cell (4, 5) yielded the best results achievable among all combinations, supporting our findings in the main paper. Furthermore, we noted that the combination of blocks 2 and 5 also achieved satisfactory reconstruction. However, employing this combination may lead to less disentanglement of style from content, as $\Delta W^{5}$ needs to “cover” $\Delta W^{2}$ by learning content details instead of focusing primarily on style, as intended. This observation further solidifies our choice of optimizing $\{\Delta W^{4},\Delta W^{5}\}$ for effective style-content separation.
@@
To validate our choice of the prompt “A [v]” during optimization, we conducted an ablation study regarding the prompt used during training. As described in the DreamBooth [45] paper, the authors suggest that the most efficient way to conduct the fine-tuning process is by using the prompt “A [v] <class-name>”, where [v] is the token dedicated for personalization, and <class-name> is the class of the object depicted in the input image. We compare our method of optimizing $\Delta W^{4}$ and $\Delta W^{5}$ with the prompt “A [v]” against using the suggested “A [v] <class-name>” prompt.
@@
Alpha Effect
As mentioned in the main paper, by the end of the training, we can obtain the tuned model weights using $W=W_{0}+\Delta W$, where $\Delta W$ is our trained B-LoRA update. The strength of the fine-tuning merge equation can be adjusted and controlled by the alpha scalar: $W=W_{0}+\alpha\Delta W$. (in our experiments $\alpha=1$). We demonstrate alpha’s effect on style and content components in Figure 22. As can be seen, when the alpha value is small, both the style and the content may be lost.
Figure 22: On the left is the style-content input pair. On the right is quantitative control over style and content by altering the $\alpha$ parameter, shown in white.
Appendix D B-LoRA for Personalization
Throughout the paper, our method has been implemented using a single image for decoupling style and content. However, by training our method using multiple images for content, we can recontextualize reference objects while preserving stylization quality. In Figures 23 and 24, we showcase the versatility of our method by combining various B-LoRAs for style and content with text prompts. Note that the style can be derived from the reference style or from other objects.
@@ Appendix D B-LoRA for Personalization
Appendix E Additional Results
Our B-LoRA method focuses on three main applications: image stylization based on image style references, text-based image stylization, and consistent style generation. In Figures 25 and 26, we present additional results generated by our approach for image stylization based on image style references. The columns represent the style reference images, while the rows correspond to the content reference images. As discussed, our method demonstrates proficiency in extracting content from style images (Figure 27) and extracting style from objects for object mixing tasks (Figure 28). In Appendices E and E, we provide qualitative results showcasing our method’s performance on randomly selected objects and styles from our evaluation set. These examples further highlight the robustness of our approach to handling diverse content and style references. In Figure 31 we present additional qualitative results for text-based image stylization. As discussed in the paper, by utilizing only the learned B-LoRA weights capturing the content, our method enables text-guided style manipulation while effectively preserving the input object’s content and structure. These results demonstrate the flexibility of our approach in allowing challenging style manipulations through textual guidance.
{NiceTabular}
@@ Appendix E Additional Results
ContentStyle   Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
@@
ContentStyle   Refer to caption Refer to caption Refer to caption Refer to caption
@@
ContentStyle   Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
@@
ContentStyle   Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
@@
{NiceTabular}
c — @   c c c c c c
[…]
Input
ContentStyle   Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
@@
{NiceTabular}
c — @   c c c c c c
[…]